OpenStudy (anonymous):

VERTICAL ASYMPTOTES? AND CHECK MY WORK y=tan(2x)-1 a. list two vertical asymptotes b. period- pi/2 c. phase shift- none d. vertical translation- down 1

OpenStudy (anonymous):

y=wat ? wat is tan

OpenStudy (anonymous):

its a trig question

OpenStudy (jdoe0001):

b. period- pi/2 \(\checkmark\) c. phase shift- none \(\checkmark\) d. vertical translation- down 1 \(\checkmark\)

OpenStudy (anonymous):

@jdoe0001 do you know how to find the vertical asymptotes, could you explain it

OpenStudy (jdoe0001):

\(\bf tan(2x)=\cfrac{sin(2x)}{cos(2x)}\qquad \textit{let's say we use }x={\color{red}{ \frac{\pi}{4}}} \\ \quad \\ tan(2x)=\cfrac{sin(2\cdot {\color{red}{ \frac{\pi}{4}}})}{cos(2\cdot {\color{red}{ \frac{\pi}{4}}})} \implies tan(2x)=\cfrac{sin(2\cdot {\color{red}{ \frac{\pi}{4}}})}{cos(\frac{\pi}{2})}\quad \textit{so, what's }cos(\frac{\pi}{2})\quad ?\)

OpenStudy (jdoe0001):

recall that for a rational, vertical asymptotes occur when the denominator turns to 0, thus making the rational "undefined" so if you just set cos(2x) = 0 and find "x" angle, those are the spots where the vertical asymptotes will occur

OpenStudy (anonymous):

im sorry that's a really good explanation im am just really really lost @jdoe0001

OpenStudy (jdoe0001):

\(\bf tan(2x)=\cfrac{sin(2x)}{cos(2x)}\qquad \textit{let's say we use }x={\color{red}{ \frac{\pi}{4}}} \\ \quad \\ tan(2x)=\cfrac{sin(2\cdot {\color{red}{ \frac{\pi}{4}}})}{cos(2\cdot {\color{red}{ \frac{\pi}{4}}})} \implies tan(2x)=\cfrac{sin(2\cdot {\color{red}{ \frac{\pi}{4}}})}{cos(\frac{\pi}{2})} \\ \quad \\ tan(2x)\cfrac{sin(2\cdot {\color{red}{ \frac{\pi}{4}}})}{0}\) so let's use set the denominator to 0 and see what we get for the angle "x" \(\bf cos(2x)=0\implies cos^{-1}[cos(2x)]=cos^{-1}(0)\implies 2x=cos^{-1}(0) \\ \quad \\ x=\cfrac{cos^{-1}(0)}{2}\Leftarrow\textit{vertical asymptotes occur at those angles}\)

OpenStudy (anonymous):

so would I plug the bottom part into my calculator?

OpenStudy (jdoe0001):

you can just look at the Unit Circle, there are 2 conspicuous spots where the cosine is 0

OpenStudy (jdoe0001):



OpenStudy (anonymous):

90 degrees and 270 degrees

OpenStudy (jdoe0001):

so that means, at those points, at those angles the cos(2x) = 0 and therefore \(\bf tan(2x)=\cfrac{sin(2x)}{cos(2x)}\) will end up with a 0 in its denominator and thus a vertical asymptote vertical asymptotes occur in a fraction when the denominator is 0

OpenStudy (anonymous):

okay how to I find those points on the graph, like not in degrees on the circle but as in the graph

OpenStudy (jdoe0001):